OpenStudy (anonymous):

Plz help Alinn flipped a coin 20 times. The coin landed heads up 12 times and tails up 8 times. Part A: Based on the results, what is the experimental probability of the coin landing heads up? Show your work. (5 points) Part B: What is the theoretical probability of the coin landing heads up? Show your work. (5 points)

OpenStudy (anonymous):

@perl

OpenStudy (perl):

experimental probability = # favorable / # total

OpenStudy (anonymous):

so 12/20 and 8/20

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply the definition of probability p, namely: \[\Large p = \frac{{favorable\;cases}}{{possible\;cases}}\]

OpenStudy (michele_laino):

now, we have 20 possible cases, and 12 favorable cases, so what is p?

OpenStudy (anonymous):

0.6

OpenStudy (michele_laino):

ok! So we can write p in percentage form, and we get 60%

OpenStudy (michele_laino):

now, in order to find the theoretical probability, we have to keep in mind that, using a coin, we have 1 favorable case and 2 possible cases, since a coin has 2 faces, namely the head face and the tail face, so what is the theoretical probability?

OpenStudy (michele_laino):

we have to apply the same formula above, namely: \[\Large p = \frac{{favorable\;cases}}{{possible\;cases}}=...?\]

OpenStudy (anonymous):

wouldn't it be the same as before with 12/20?

OpenStudy (michele_laino):

no, since that is an experimental probability

OpenStudy (anonymous):

ok so I'm kinda confused with the theoretical part

OpenStudy (michele_laino):

the theoretical probability, refers to the coin. Namely a coin has 2 faces, head and tail. So there are two possible outcomes: head or tail. Now we want a head so we have 2 possible cases and 1 favorable case

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

whereas the experimental probability refers to the experiment that we have made. In our case we have got 12 heads (favorable cases) in 20 times (possible cases), and, as we have computed before, the probability is: 12/20

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

so, what is the theoretical probability? keep in mind that we have 1 favorable case ( a head) and 2 possible cases (the number of the face of a coin)

OpenStudy (michele_laino):

faces*

OpenStudy (michele_laino):

\[\Large p = \frac{{favorable\;cases}}{{possible\;cases}} = ...?\]

OpenStudy (anonymous):

0.5

OpenStudy (michele_laino):

that's right! As usual, we can rewrite that probability in percentage form, and we get: p=50%

OpenStudy (anonymous):

ok